# Diversity as a Reward: Fine-Tuning LLMs on a Mixture of Domain-Undetermined Data

### v2

@@ Abstract
11footnotetext: Equal contribution.
@@ 1 Introduction
To cultivate comprehensive capabilities in LLMs, prior studies have explored preferable trade-offs between quality, quantity, and diversity of their training data  li2024quantity ; djv2 ; qin2024surveycodev ; zhao2024beyond . For example, methods focusing on data selection li2024super ; xialess ; wang2023far and mixture ge2024bimix demonstrate promising capabilities to enhance model performance, particularly through semantic diversity  lu2023instag ; liu2023deita .
@@
These findings motivate our theoretical analysis, which establishes that traditional data mixing strategies assuming known domain weights ye2024data cannot achieve Pareto-optimal performance in label-free settings. We prove that when domain labels are unavailable, the optimal data selection strategy must simultaneously maximize the mutual information between selected samples and preserve the base model’s latent feature geometry.
@@
Extensive evaluations across 7 benchmarks and 3 model families reveal that DaaR achieves new state-of-the-art (SOTA) average performance, consistently outperforming 9 baseline methods on high-difficulty tasks while maintaining computational efficiency. Crucially, our method requires no domain labels or external models - the LLM self-supervises its diversity estimation through geometric constraints in its native embedding space.
Our core contributions for LLM data optimization are three-fold:
@@
  • A new method with label-free diversity reward mechanism using embedding-space entropy prediction, theoretically grounded in mutual information maximization.
@@ 2.1 Related Works
Fine-tuning is a pivotal training paradigm for enhancing LLMs’ domain-specific capabilities. Research has shown that a small set of instruction pairs can enable LLMs to follow major instructions effectively zhou2024lima ; djv1 . This fine-tuning can be achieved through rule-based methods, which focus on attributes such as Error L2-Norm paul2021el2n and token length raffel2020exploring . More recently, model-based heuristics have been explored, including methods based on instruction-following difficulty li2024super , GPT scoring chen2024alpagasus ; 2024qurating , data model selection 2024dsdm , and influence scores derived from loss xialess .
@@
A fundamental principle for LLMs is being able to handle diverse human requests, underscoring data diversity as essential for effective fine-tuning yu2022can ; ding2023enhancing . Data diversity encompasses aspects such as data deduplication abbas2023semdedup , the coverage scope of tags lu2023instag , model-based diversity evaluation liu2023deita ; zhang2024harnessing , and scaling properties song2024scaling . Typically, data mixing approaches albalak2023efficient ; ye2024data ; ge2024bimix focus on adjusting the proportional weights of different domains to enhance model capabilities.
@@ 3.1 Seed Data Pools and Basic Setting
All experiments are conducted under identical training and evaluation protocols with two independent repetitions. Full platform, training and evaluation details are provided in Appendix B.
@@ 3.3 Experimental Observations
Table 1 presents comprehensive evaluations across seven benchmarks, where the notation Inter-Diversity (X-Y) indicates samples ranked in the top (100-Y)to (100-X)of cross-domain similarity scores. Due to space constraints, we present only the results for the top 20%, middle 20%, and bottom 20%. Detailed performance results on various downstream tasks are shown in Appendix H.4. Our diversity-controlled selections reveal two observations:
@@ 3.4 Theoretical Perspective
[…]
@@
The derivation is detailed in Appendix E. This decomposition reveals that the Inter-Diversity and Intra-Diversity are governed by $\lambda_{k}$ and $p(c{=}k|x)$, leading to the following analysis.
@@
This formulation is commonly used when domain labels are available, under the implicit assumption that domains are mutually exclusive. However, such an assumption rarely holds in the context of LLM training, as cross-domain data may either synergize or conflict zhao2024beyond ; xie2023doremi (evidence in Table 1). Hence, neglecting diversity in data selection potentially leads to suboptimal overall performance.
@@
Under the domain-undetermined scenario, our proposed method DaaR operationalizes the gained insights by (1) constructing pseudo-labels (Section 4.1) align with LLM-aware generated seed, (2) explicitly modeling the classifier $p(c{=}k|x)$ through a domain discrimination probe (Section 4.2), and (3) using the probe’s predictive entropy (Section 4.2) as a proxy for diversity, estimating the ability of classifier $p(c{=}k|x)$ to select data toward the target distribution $q$.
@@ 4 DaaR: Diversity as a Reward
To address the challenges identified in Sec. 3.3 and leverage the insights gained in Sec. 3.4, we establish a data selection method DaaR guided by diversity-aware reward signals. It comprises three key components illustrated in Fig. 1: (1) model-aware centroid synthesis, (2) two-stage training with reward probe, and (3) diversity-driven data selection.
@@ 4.1 Model-Aware Training Data
  • Phase 1 - Seed Generation: For each domain $k$, generate seed samples $\mathcal{S}_{k}^{(0)}$ via zero-shot prompting with domain-specific description templates generated by LLM itself, alongside with minor injection from downstream task samples, establishing initial semantic anchors.
  • Phase 2 - Diversity Augmentation: Iteratively expand $S_{k}^{(t)}$ through context-aware generation, conditioned on a sliding window buffer with 3 random anchors sampled from the $(t-1)$ iteration $S_{k}^{(t-1)}$. The generated sample $\mathbf{x^{\prime}}$ is retained through rejection sampling to enhance diversity.:
@@
    where $\mathcal{M}_{\text{ebd}}(\cdot)$ indicates the output of the embedding layer of the given LLM, with $\tau$ as the similarity threshold. This process terminates when it reaches the predetermined iteration. The implementation of all parameters, along with their corresponding ablation studies, is provided in Appendix C.1.
@@ 4.2 Training for Self-Rewarding Abilities
The proposed DaaR establishes model-aware domain discrimination abilities through a multi-layer perceptron (MLP) probe module, $\psi_{\text{dom}}$, attached to the hidden layer $\mathcal{M}(\tilde{x})$ of the LLMs. The probe will be trained meanwhile all the parameters of the LLM are frozen, achieving a preferable balance between effectiveness and cost, with detailed analysis regarding the choice of layers presented in Appendix G.2. Specifically, with pseudo-label $\tilde{y}$, we can compute domain predicted probabilities as:
@@
$\psi_{\text{dom}}$ is optimized via cross-entropy loss $\mathcal{L}_{\text{dom}}=-\frac{1}{|\mathcal{D}_{\text{probe}}|}\sum_{(\tilde{x},\tilde{y})\in\mathcal{D}_{\text{probe}}}\sum_{k=1}^{K}\mathbb{I}_{[k=\tilde{y}]}\log p_{k}(\tilde{x})$, where $\mathbb{I}_{[k=\tilde{y}]}$ denotes the indicator function. We employ single-sample batches with the AdamW optimizer to prevent gradient averaging across domains. Training consistently converges and achieves 92.7% validation accuracy on domain classification, as shown in Fig. 2 (a).
@@
Building on the stabilized domain discrimination probe, we quantify sample-level diversity through predictive entropy $H(\tilde{x})$. And to enable efficient reward computation during data selection, we then train another 5-layer MLP $\psi_{\text{div}}$ to directly estimate $H(\tilde{x})$ from $\mathcal{M}(\tilde{x})$:
@@ 5 Experiments
We use the following data selection methods for comprehensive evaluation: (1) Random Selection: traditional random sampling; (2) Instruction Len: measuring instruction complexity by token count  cao2023instruction ; (3) Alpagasus chen2024alpagasus : using ChatGPT for direct quality scoring of instruction pairs; (4-5) Instag lu2023instag : semantic analysis approach with Instag-C (complexity scoring via tag quantity) and Instag-D (diversity measurement through tag set expansion); (6) SuperFilter li2024super : response-loss-based complexity estimation using compact models; (7-9) Deita liu2023deita : model-driven evaluation with Deita-C (complexity scoring), Deita-Q (quality scoring), and Deita-D (diversity-aware selection). Detailed implementations for these baselines are in Appendix D.1.
@@ 5.1 Overall Performance
: The task of balanced capability enhancement proves particularly challenging for existing methods. While some baselines achieve strong performance on specific tasks (e.g., SuperFilter’s 40.55 on HumanEval for Llama3.1), they suffer from catastrophic performance drops in other domains (e.g., SuperFilter’s 6.05 on MATH). Only 3 baselines perform better than Random on Llama3.1-8B. In the Qwen series, even all baselines fall below the Random performance. We conjecture this stems from selected distribution-bias, as the baselines are unable to balance the proportions across different domains, a perspective visually supported and analyzed in Appendix D.2.
Comprehensive Performance
5.2 Generalizability and Uniqueness of DaaR
Evaluation on MMLU In the main experiments, the data and benchmarks we employed are meticulously aligned, which validates the In-Distribution capabilities of DaaR. To further validate the generalization ability of DaaR, we present MMLU mmlu2021 evaluations in Appendix F.1. MMLU encompasses a significantly broader range of content than our dataset, thus serving as an Out-of-Distribution scenario. Results demonstrate that DaaR achieved top accuracy on Qwen2 with an improvement of 0.61 points, and secured a competitive second-place performance on Qwen2.5. At a granular level, DaaR particularly excels in the STEM domain, outperforming baselines by an average of 7.6% and 1.7% on Qwen2 and Qwen2.5, which aligns the findings above.
Evaluation on Qwen3 In Appendix F.2, we explore the performance of DaaR on the mainstream RL-based LLM, Qwen3-8B. We observe that despite difference in structure, the phenomena (Fig. 8 & Table 8) and the dynamics of training (Fig. 9) remain consistent with the main observations. Then we compare DaaR with baselines in Table 9, although Qwen3-8B is not directly designed for SFT paradigm, DaaR maintains the top position in Avg performance across baselines, surpasses all baselines up to 4.7% and notably exceeds the avg of baselines by 25.91% on MATH.
Customized Characteristics Unlike baselines, DaaR enables domain-aware data selection in undetermined scenario, showcasing potential customization on $\lambda_{k}$. As detailed in Appendix F.3, by adjusting the selection ratio for specific domains, we achieve SOTA performance in targeted domain without compromising overall ability, promising practical utility in real-world application.
Table 3: Illustration results on MMLU, Qwen3 and customization, detailed in Table 6, 9, 10
@@ 5.3 Ablation Studies and Discussions
Distinct Domain Centroid As shown in Appendix G.1, DaaR can generate domain-representative samples with clear distinctions. Notably, the diversity is consistently most pronounced between common sense, reasoning, and coding domains across model architectures and parameters.
Impacts of Hyper-Parameters In Appendix G.2-Appendix G.6, we conduct a series of ablation studies, including (1) selection of layer: the discrimination accuracy derived from layer 1-10 fluctuates in 91.2-93.6%; (2) the number of seed: the similarity between centroids remains highly consistent with variance ranging from 5.4E-7 to 1.8E-5; (3) the number of augmented data: the mean similarity of centroids generated from sizes of 10 and 30 is 0.983 and 0.981 on Qwen2.5 and Llama3.1 respectively; (4) the size of sliding window & similarity threshold: the impact of different hyper-parameters on centroid generation is minimal, ensuring the subsequent training stability of DaaR.
Cost-Efficiency and Flexibility As shown in Appendix G.7, our method demonstrates significant computational efficiency compared to baselines that rely on GPT-based evaluators or full-LLM inference. By operating on frozen hidden layers and generating regression scores instead of full vocabulary distributions, our approach achieves 70% lower GPU usage and 2.5x faster inference.
6 Conclusions, Limitations & Future Works

### v3

@@ Abstract
11footnotetext: Equal contribution.22footnotetext: Corresponding author.33footnotetext: FSIETP: Guangdong Provincial Key Laboratory of Fire Science and Intelligent Emergency Technology.
@@ 1 Introduction
To cultivate comprehensive capabilities in LLMs, prior studies have explored preferable trade-offs between quality, quantity, and diversity of their training data  li2024quantity ; djv1 ; qin2024surveycodev ; zhao2024beyond . For example, methods focusing on data selection li2024super ; xialess ; wang2023far and mixture ge2024bimix demonstrate promising capabilities to enhance model performance, particularly through semantic diversity  lu2023instag ; liu2023deita .
@@
These findings motivate our theoretical analysis, which establishes that traditional data mixing strategies assuming known domain weights ye2024data cannot achieve Pareto-optimal performance in label-free settings. Our analysis further reveals that when domain labels are unavailable, an effective data selection strategy should prioritize samples that enhance sample-level diversity. This strategy involves balancing the goal of enhancing sample-level diversity to approximate the ideal importance weights, with the need to preserve the base model’s latent feature geometry.
@@
Extensive evaluations across 7 benchmarks and 3 model families reveal that DaaR achieves new state-of-the-art (SOTA) average performance, consistently outperforming 9 baseline methods on high-difficulty tasks while maintaining computational efficiency. Crucially, our method requires no domain labels or external models - the LLM self-supervises its diversity estimation through geometric constraints in its native embedding space. Our core contributions for LLM data optimization are:
@@
  • A new method with label-free diversity reward mechanism using embedding-space entropy prediction, theoretically grounded in importance sampling.
@@ 2.1 Related Works
Fine-tuning is a pivotal training paradigm for enhancing LLMs’ domain-specific capabilities. Research has shown that a small set of instruction pairs can enable LLMs to follow major instructions effectively zhou2024lima ; mindgym . This fine-tuning can be achieved through rule-based methods, which focus on attributes such as Error L2-Norm paul2021el2n and token length raffel2020exploring . More recently, model-based heuristics have been explored, including methods based on instruction-following difficulty li2024super , GPT scoring chen2024alpagasus ; 2024qurating , data model selection 2024dsdm , and influence scores derived from loss xialess .
@@
A fundamental principle for LLMs is being able to handle diverse human requests, underscoring data diversity as essential for effective pre-training liu2024regmix ; chen2024aioli ; xie2025chameleon ; fan2024dynamic and fine-tuning yu2022can ; ding2023enhancing . Data diversity encompasses aspects such as data deduplication abbas2023semdedup , the coverage scope of tags lu2023instag , model-based diversity evaluation liu2023deita ; zhang2024harnessing , and scaling properties song2024scaling . Typically, data mixing approaches albalak2023efficient ; ye2024data ; ge2024bimix ; liu2024regmix focus on adjusting the proportional weights of different domains to enhance model capabilities.
@@ 3.1 Seed Data Pools and Basic Setting
All experiments are conducted under identical training and evaluation protocols with two independent repetitions. Full platform, training and evaluation details are provided in Appendix D.
@@ 3.3 Experimental Observations
Table 1 presents comprehensive evaluations across seven benchmarks, where the notation Inter-Diversity (X-Y) indicates samples ranked in the top (100-Y)% to (100-X)% of cross-domain similarity scores. Due to space constraints, we present only the results for the top 20%, middle 20%, and bottom 20%. Detailed performance results on various downstream tasks are shown in Appendix H.4. Our diversity-controlled selections reveal two observations:
@@ 3.4 Theoretical Perspective
$\small\mathbb{E}_{x\sim q(\cdot)}\ell(x)=\int_{x}q(x)\ell_{\theta}(x)dx=\int_{x}p(x)\frac{q(x)}{p(x)}\ell_{\theta}(x)dx=\mathbb{E}_{x\sim p(\cdot)}\frac{q(x)}{p(x)}\ell(x;\theta)=\sum_{x\sim\mathcal{D}}\frac{q(x)}{p(x)}\ell(x;\theta).$ (4)
@@
The derivation is detailed in Appendix C.1. This decomposition reveals that the Inter-Diversity and Intra-Diversity are governed by $\lambda_{k}$ and p(c=k|x)p(c{=}k|x), leading to the following analysis.
@@
This formulation is commonly used when domain labels are available, under the implicit assumption that domains are mutually exclusive. However, such an assumption rarely holds in the context of LLM training, as cross-domain data may either synergize or conflict zhao2024beyond ; xie2023doremi ; kuang2025atomic (evidence in Table 1). Hence, neglecting diversity in data selection potentially leads to suboptimal overall performance. We further formalize this suboptimality by analyzing the approximation error incurred by such a deterministic assignment, which provides a supplementary justification for using predictive entropy as a selection criterion. For a detailed analysis, please refer to Appendix C.2.
@@
Under the domain-undetermined scenario, our proposed method DaaR operationalizes the gained insights by (1) constructing pseudo-labels (Section 4.1) align with LLM-aware generated seed, (2) explicitly modeling the classifier p(c=k|x)p(c{=}k|x) through a domain discrimination probe (Section 4.2), and (3) using the probe’s predictive entropy (Section 4.2) as a proxy for diversity, estimating the ability of classifier p(c=k|x)p(c{=}k|x) to select data toward the target distribution $q$.
@@ 4 DaaR: Diversity as a Reward
To address the challenges identified in Sec. 3.3 and leverage the insights gained in Sec. 3.4, we establish a data selection method DaaR guided by diversity-aware reward signals.
It comprises three key components illustrated in Fig. 1: (1) model-aware centroid synthesis, which generates domain-representative centroids and seed data capturing the LLM’s intrinsic feature space, (2) two-stage training with reward probe, which yields a probe module capable of predicting sample-level diversity rewards accurately, and (3) diversity-driven data selection to obtain a data subset that effectively boosts LLM’s cross-domain performance.
@@ 4.1 Model-Aware Training Data
  • Phase 1 - Seed Generation: For each domain $k$, generate seed samples $\mathcal{S}_{k}^{(0)}$ via zero-shot prompting with domain-specific description templates generated by LLM itself, alongside with minor injection from downstream task samples, establishing initial semantic anchors. We show that removing the minor injection of downstream samples negligibly impacts final performance but significantly reduces data generation efficiency in Appendix G.8.
  • Phase 2 - Diversity Augmentation: Iteratively expand $S_{k}^{(t)}$ through context-aware generation, conditioned on a sliding window buffer with 3 random anchors sampled from the (t1)(t-1) iteration $S_{k}^{(t-1)}$. The generated sample $\mathbf{x^{\prime}}$ is retained through rejection sampling to enhance diversity.:
@@
    where $\mathcal{M}_{\text{ebd}}(\cdot)$ indicates the output of the embedding layer of the given LLM, with $\tau$ as the similarity threshold. This process terminates when it reaches the predetermined iteration.
@@
More implementation details including hyper-parameters and the data generation process, along with their corresponding ablation studies, are provided in Appendix B.
@@ 4.2 Training for Self-Rewarding Abilities
[…]
@@
[…]
@@
[…]
@@ 5 Experiments
We use the following data selection methods for comprehensive evaluation: (1) Random Selection: traditional random sampling; (2) Instruction Len: measuring instruction complexity by token count  cao2023instruction ; (3) Alpagasus chen2024alpagasus : using ChatGPT for direct quality scoring of instruction pairs; (4-5) Instag lu2023instag : semantic analysis approach with Instag-C (complexity scoring via tag quantity) and Instag-D (diversity measurement through tag set expansion); (6) SuperFilter li2024super : response-loss-based complexity estimation using compact models; (7-9) Deita liu2023deita : model-driven evaluation with Deita-C (complexity scoring), Deita-Q (quality scoring), and Deita-D (diversity-aware selection). Detailed implementations for these baselines are in Appendix E.1.
@@ 5.1 Overall Performance
: The task of balanced capability enhancement proves particularly challenging for existing methods. While some baselines achieve strong performance on specific tasks (e.g., SuperFilter’s 40.55 on HumanEval for Llama3.1), they suffer from catastrophic performance drops in other domains (e.g., SuperFilter’s 6.05 on MATH). Only 3 baselines perform better than Random on Llama3.1-8B. In the Qwen series, even all baselines fall below the Random performance. We conjecture this stems from selected distribution-bias, as the baselines are unable to balance the proportions across different domains, a perspective visually supported and analyzed in Appendix E.2.
Table 3: Illustration results on MMLU, Qwen3 and customization, detailed in Table 7, 10, 11
@@
Comprehensive Performance
5.2 Generalizability, Uniqueness and Stability of DaaR
Evaluation on MMLU
In the main experiments, the data and benchmarks we employed are meticulously aligned, which validates the In-Distribution capabilities of DaaR. To further validate the generalization ability of DaaR, we present MMLU mmlu2021 evaluations in Appendix F.1. MMLU encompasses a significantly broader range of content than our dataset, thus serving as an Out-of-Distribution scenario. Results demonstrate that DaaR achieved top accuracy on Qwen2 with an improvement of 0.61 points, and secured a competitive second-place performance on Qwen2.5. At a granular level, DaaR particularly excels in the STEM domain, outperforming baselines by an average of 7.6% and 1.7% on Qwen2 and Qwen2.5, which aligns the findings above.
Evaluation on Qwen3
In Appendix F.2, we explore the performance of DaaR on the mainstream RL-based LLM, Qwen3-8B. We observe that despite difference in structure, the phenomena (Fig. 9 & Table 9) and the dynamics of training (Fig. 10) remain consistent with the main observations. Then we compare DaaR with baselines in Table 10, although Qwen3-8B is not directly designed for SFT paradigm, DaaR maintains the top position in Avg performance across baselines, surpasses all baselines up to 4.7% and notably exceeds the avg of baselines by 25.91% on MATH.
Customized Characteristics
Unlike baselines, DaaR enables domain-aware data selection in undetermined scenario, showcasing potential customization on $\lambda_{k}$. As detailed in Appendix F.3, by adjusting the selection ratio for specific domains, we achieve SOTA performance in targeted domain without compromising overall ability, promising practical utility in real-world application.
Table 4: DaaR using model-aware pseudo-labels vs. Ground-Truth (GT) labels.
Model Setting NQ TriviaQA Hellaswag GSM8k MATH MBPP HumanEval Avg
Llama3.1-8B DaaR w/ Pseudo 20.08 64.55 74.88 54.80 15.30 4.70 37.50 38.83
DaaR w/ GT 22.54 65.52 73.43 53.80 14.35 4.40 36.50 38.65
Diff +2.46 +0.97 -1.45 -1.00 -0.95 -0.30 -1.00 -0.18
Qwen2-7B DaaR w/ Pseudo 16.88 57.58 73.03 75.40 38.10 52.00 64.94 53.99
DaaR w/ GT 18.75 58.35 72.02 73.40 35.75 51.50 64.01 53.40
Diff +1.87 +0.77 -1.01 -2.00 -2.35 -0.50 -0.93 -0.59
Qwen2.5-7B DaaR w/ Pseudo 15.83 58.65 72.48 80.20 16.70 64.20 68.29 53.76
DaaR w/ GT 16.04 58.75 71.87 79.40 16.30 64.25 67.76 53.48
Diff +0.21 +0.10 -0.61 -0.80 -0.40 +0.05 -0.53 -0.28
Robustness and Stability
We validate our model-aware design choices that relying on external ground-truth labels can degrade performance by up to 0.59 points in Table 4 and Appendix F.4. Furthermore, we demonstrate high end-to-end robustness, with independent runs yielding inter-centroid similarities of >0.98 and a final performance standard deviation of only 0.08 in Appendix F.5
@@ 5.3 Ablation Studies and Discussions
Ablation Studies on Centroid Construction
As shown in Fig.3 and Appendix G.1, DaaR can generate domain-representative samples with clear distinctions. Notably, the diversity is consistently most pronounced between common sense, reasoning, and coding domains across model architectures and parameters, showing that despite differences in model architecture across LLMs, their discriminability between these domains exhibit remarkable similarity.
Impacts of Hyper-Parameters
In Appendix G.2-Appendix G.7, we conduct a series of ablation studies, including (1) selection of layer: the discrimination accuracy derived from layer 1-10 fluctuates in 91.2-93.6%; (2) the number of seed: the similarity between centroids remains highly consistent with variance ranging from 5.4E-7 to 1.8E-5; (3) the number of augmented data: the mean similarity of centroids generated from sizes of 10 and 30 is 0.983 and 0.981 on Qwen2.5 and Llama3.1 respectively; (4) the size of sliding window & similarity threshold: the impact of different hyper-parameters on centroid generation is minimal, ensuring the subsequent training stability of DaaR; (5) diversity threshold, where varying thresholds minimally affect centroid-guided clustering (with distinct generation attempt frequencies across thresholds) and empirically balanced thresholds are determined for different domains; (6) generalizability across model scales, where DaaR shows consistent improvements in representative scenario on both Llama3.2-3B and Qwen2.5-14B models.
Cost-Efficiency and Flexibility
As detailed in Appendix G.9, our method demonstrates significant computational efficiency compared to baselines that rely on GPT-based evaluators or full-LLM inference. By operating on frozen hidden layers and generating regression scores instead of full vocabulary, our approach achieves 70% lower GPU usage and 2.5x faster inference.
6 Conclusion, Limitation & Future Works